In a E.coli 4.6*106 bps are present then how many nucleosome are present in this 
1) 2.3 * 104
2) 2.3 * 108
3) 1.65 * 107
4) zero

Grade:12th pass

5 Answers

Simran Vinaik
5 years ago
Ans is zero as Ecoli is a prokaryote so its doesnt have histone proteins/ nucleosome
Shivi .
5 years ago
What if it was not a prokaryotic organism and a eukaryotic then how can we solve such problems .what is the method ?
Simran Vinaik
5 years ago
One nucleosome has 147 bp of DNA. So just divide the total bp by 147 to get no of nucleosome
Varsha RC
4 years ago
ecoli has 4.6*106 bps 
nucleosome has 200 bps
here 1 nucleosome = 200bp
         x nucleosome= 4.6*106 bps
      
      x =4.6*106 bps/200 = 2.3 * 10
 
 ans)1
Chailsi Mittal
3 years ago
zero nucleosome as Ecoli is an prokaryote and it does not have histone protein instead of it have polyamines ..so answer will be 4th option